# PICASSO: A Feed-Forward Framework for Parametric Inference of CAD Sketches via Rendering Self-Supervision

### v1

@@ Abstract
We propose PICASSO, a novel framework CAD sketch parameterization from hand-drawn or precise sketch images via rendering self-supervision. Given a drawing of a CAD sketch, the proposed framework turns it into parametric primitives that can be imported into CAD software. Compared to existing methods, PICASSO enables the learning of parametric CAD sketches from either precise or hand-drawn sketch images, even in cases where annotations at the parameter level are scarce or unavailable. This is achieved by leveraging the geometric characteristics of sketches as a learning cue to pre-train a CAD parameterization network. Specifically, PICASSO comprises two primary components: (1) a Sketch Parameterization Network (SPN) that predicts a series of parametric primitives from CAD sketch images, and (2) a Sketch Rendering Network (SRN) that renders parametric CAD sketches in a differentiable manner. SRN facilitates the computation of an image-to-image loss, which can be utilized to pre-train SPN, thereby enabling zero- and few-shot learning scenarios for the parameterization of hand-drawn sketches. Extensive evaluation on the widely used SketchGraphs dataset [36] validates the effectiveness of the proposed framework.
@@ 1 Introduction
Computer-Aided Design (CAD) has become the industry norm for mechanical design of any product prior to manufacturing. CAD software [3, 2] enhances the productivity of engineers and enables efficient extension or alteration of existing designs. The modern CAD workflow widely adopts the paradigm of feature-based modeling [43], where initially a series of two-dimensional parametric CAD sketches is specified, followed by CAD operations (e.g. extrusion, revolution, etc.) to form a 3D solid. CAD sketches comprise a collection of geometric primitives (e.g. lines, arcs, points) as well as geometric constraints enforced between those primitives (e.g. coincident, parallel, etc.). Feature-based CAD constitutes an efficient way of constructing complex 3D models [19, 11, 25, 12]. Commonly, the modelling process starts with conceptualization of a design by roughly drawing it by hand [37]. Designers are tasked with meticulously translating these drawings, often in the form of raster images, into parametric CAD sketches. Depending on the complexity of the target design, the task can be time-consuming even for highly skilled designers. Thus, the automation of the CAD parameterization process has gained attention within the research community [45, 37] and CAD industry [1, 3].
Figure 1: PICASSO enables zero- and few-shot learning of CAD sketch parameterization from hand-drawn raster sketches.
@@
Figure 2: Framework overview. PICASSO is composed of two networks, namely the Sketch Parametrization Network (SPN) and Sketch Rendering Network (SRN). Once trained, SRN is kept frozen and used for rendering self-supervision using a multiscale loss. This allows for image-level pre-training of the CAD sketch parameterization network SPN.
@@
Parameterization of CAD sketches from raster images constitutes a complex problem due to the large solution space required to model parametric entities, the nuanced intricacies of sketch designs, and the inherently inaccurate nature of hand-drawings. Compared to vectorized stroke inputs [4], raster images pose additional challenges due to indistinguishable, overlapping or closely positioned primitives on the image domain. The availability of large-scale CAD sketch datasets [36, 14] has enabled tackling this problem with learning-based approaches. Recent works [14, 28, 37] propose autoregressive generative models to learn CAD sketch parameterization through general-purpose language modelling. Similarly to next word prediction for natural language processing (NLP), these models predict the next primitive in the sketch and can be conditioned on images for sketch parameterization. Transformer-based networks are typically trained to predict tokens representing the type and parameters of each primitive via parameter-level supervision. These networks are adept at exploiting large-scale annotated datasets with millions of sketches (e.g. SketchGraphs [36]). Nevertheless, there are real-world situations where this quantity of labeled sketches is unavailable. Examples include rough hand-drawn CAD sketches and 2D cross-sections [6] from 3D scans, where collecting parameter annotations is challenging. These examples exhibit a distribution shift w.r.t. to CAD sketches in large-scale datasets, underscoring the necessity of learning CAD sketch parametrization from raster images when parameter-level annotations are scarce or nonexistent.
@@
In this work, we introduce a framework for Parametric Inference of CAD Sketches via Rendering Self-SupervisiOn, refereed to as PICASSO. PICASSO enables learning parametric CAD sketches directly from precise or hand-drawn images, even when parameter-level annotations are limited or unavailable. This is achieved by utilizing the geometric appearance of sketches as a learning signal to pretrain a CAD parameterization network. Specifically, PICASSO is composed of two main components: (1) a Sketch Parameterization Network (SPN) that predicts a set of parametric primitives from CAD sketch images and (2) a Sketch Rendering Network (SRN) to render parametric CAD sketches in a differentiable manner. SRN enables image-to-image loss computation that can be used to pretrain SPN, leading to zero- and few-shot learning scenarios for hand-drawn sketch parametrization (see Figure 1). To the best of our knowledge, we are the first to address CAD sketch parameterization with limited or without parametric annotations. PICASSO can achieve strong parameterization performance with only a small number of annotated samples. Moreover, departing from recent work based on autoregressive modelling [37, 28, 14], the proposed SPN is a feed-forward network. The sketch is treated as an unordered set of primitives and a CAD sequence is predicted in a non-autoregressive manner. Experiments show that SPN outperforms recent autoregressive state-of-the-art on parameterizing precise and hand-drawn CAD sketch images. An overview of PICASSO is presented in Figure 2.
@@ 2 Related Works
Concurrently, SketchGen [28] and Vitruvion [37] propose two-stage architectures for both primitive and constraint generation. Seff et al. [37] also investigate their model’s efficacy for the parameterization of hand-drawn sketches. These methods follow the autoregressive learning strategy as it constitutes the natural choice for generative modelling. Autoregressive inference is also well-suited for the related application of sketch auto-completion but might be suboptimal for CAD parameterization from images (see detailed discussion in Section 4.2). In contrast, the proposed SPN is a feed-forward network. Related to ours is also the non-autoregressive method of [45]. Authors introduce the task of modular concept discovery that is addressed through a program library induction perspective. The proposed framework discovers a modular library and reconstructs parametric sketches by module composition and instantiation of module parameters. Recently, the authors in [39] presented a concurrent work to ours. They highlighted the issue of error accumulation in autoregressive models and introduced a feed-forward strategy for CAD sketch parameterization, which is similar to our Sketch Parameterization Network (SPN). All aforementioned methods solely rely on parameter-level supervision and overlook the rendered geometry of a CAD sketch. In this work, rendering self-supervision is proposed to provide an alternative signal for sketch parameterization. Such signal allows for pre-training CAD sketch parameterization network, hence enabling it to parameterize hand-drawn or precise CAD sketches even under limited or without parametric supervision.
@@ 3 Problem Statement
Figure 3: Overview of the Sketch Rendering Network (SRN). SRN is modeled by a transformer encoder-decoder that learns a mapping from parametric primitive tokens to the sketch image domain. Through neural differentiable rendering, SRN allows the computation of an image-to-image loss between predicted raster sketches and input precise or hand-drawn sketches.
Our goal becomes to learn the mapping $\boldsymbol{F}_{\theta}:\{0,1\}^{h\times w}\rightarrow\mathcal{P}^{n}$ with model parameters $\theta$, from the image domain to the space of parametric primitives. As in [37], each primitive is represented by a set of tokens defining primitive types and quantized primitive parameters. A $6$-bit uniform quantization is adopted. More specifically, every primitive is expressed as a set of 8 tokens $\mathbf{p}_{i}~{}=~{}\{t_{i}^{j}\}_{j\in\llbracket 1..8\rrbracket}$ with $t_{i}^{j}\in\llbracket 0..72\rrbracket$. Tokens in $\llbracket 0..6\rrbracket$ represent primitive types i.e., padding, start, end, arc, circle, line, and point, respectively. Tokens in the range $\llbracket 7..70\rrbracket$ correspond to quantized primitive parameters and values $\{71,72\}$ refer to whether the primitive is a construction primitive (used by designers for referencing). The complete set of quantized primitives can be formally defined as ${\mathbf{y}}~{}\in~{}\{0,1\}^{8n\times 73}$ where $8$ represents the tokens for each primitive and $73$ is the tokenization interval.
@@ 4.1 Sketch Rendering Network
This work proposes a Sketch Rendering Network (SRN) designed for neural rendering of parametric CAD sketches. SRN learns $\mathbf{\Phi}_{\phi}:~{}\mathcal{P}^{n}~{}\rightarrow~{}\{0,1\}^{h\times w}$, i.e., the mapping from a set of parametric token primitives to sketch image domain. The primary utility of SRN is to allow the computation of an image-to-image loss between predicted primitives and their corresponding precise or hand-drawn input sketch image, thus enabling image-level pre-training for CAD sketch parametrization. Note that explicit rendering of CAD sketches implies the direct rasterization of parametric primitives, and it is inherently a non-differentiable process. While differentiable rendering solutions exist in literature [22], their use is limited to parameter refinement.
Network Architecture: The proposed architecture for SRN is depicted in Figure 3. SRN operates on one-hot encoded tokens ${\mathbf{y}}~{}\in~{}\{0,1\}^{8n\times 73}$. Since input tokens correspond to an unordered set of parametric primitives (as mentioned in Section 3), SRN adopts an encoder-decoder transformer architecture. Firstly, the tokens are projected through a linear layer $\mathbf{\Phi}_{f}$ to produce token embedding $\mathbf{f}_{r}~{}\in~{}\mathbb{R}^{8n\times d_{q}}$, where $d_{q}$ denotes the embedding dimension. The formed embeddings are processed by a transformer encoder network $\mathbf{\Phi}_{e}$ without changing their dimensions. This is followed by a decoder $\mathbf{\Phi}_{d}$, inspired by [16], that maps via cross-attention a set of randomly initialized patch queries $\mathbf{q}_{\Phi}~{}\in~{}\mathbb{R}^{\frac{hw}{d_{e}^{2}}\times d_{e}\times d_{e}}$ to a set of decoded patches of the same dimension. Here, $d_{e}$ denotes the patch size. The final rendered image is formed by rearranging those patches and applying a linear layer with sigmoid activation.
@@ 4.2 Sketch Parametrization Network
Network Architecture: Figure 4 shows the architecture of SPN. First, a convolutional [34] backbone $\mathbf{\Psi}_{f}$ extracts a set of image features $\mathbf{f}\in\mathbb{R}^{c\times h\times w}$, for input sketch image $\mathbf{X}$. The input feature map is divided into non-overlapping patches that are embedded along with fixed positional embeddings [29] and passed through a standard vision transformer encoder [10] $\mathbf{\Psi}_{v}$ to produce a set patches $\mathbf{v}~{}\in~{}\mathbb{R}^{\frac{hw}{d_{e}^{2}}\times d_{e}\times d_{e}}$. The decoder $\mathbf{\Psi}_{d}$ is a vanilla transformer [38] trained to decode a set of $d$ fixed query vectors $\mathbf{q}_{\Psi}~{}\in~{}~{}\mathbb{R}^{d\times d_{q}}$ into primitive tokens via self and cross-attention. Note that $d_{e}$ and $d_{q}$ maintain the same value as for SRN. The decoded tokens are classified independently via a linear layer with softmax activation resulting in token probabilities $\hat{\mathbf{y}}~{}\in~{}[0,1]^{8n\times 73}$.
@@ 4.3 Rendering Self-Supervision for CAD Sketch Parameterization
$\mathcal{L}_{ml2}=\sum_{s\in S}\|d_{s}\left(\mathbf{\Phi}_{\phi}(\boldsymbol{F}_{\theta}(\mathbf{X})\right))-d_{s}\left(\mathbf{X}\right)\|_{2}^{2}\ ,$ (1)
@@ 5.1 Experimental Setup
Implementation Details: Images are of size $h~{}=~{}128$ and $w=128$. SPN uses a U-Net [34] convolutional backbone with a ResNet34 [17] that produces a $16\times 128\times 128$ feature map. For SPN, the transformer encoder-decoder is formed by $4$ layers each, with $8$ heads and $d_{q}=256$ latent dimensions. The SRN transformer has $12$ layers with $8$ heads and the same latent dimension as SPN. Patch size $d_{e}$ is set to $16$. As in [37], the maximum number of primitives is fixed to $n=16$. We train SRN and SPN for $40$ and $20$ epochs respectively, with a learning rate of $1\times 10^{-4}$ and a batch size of $128$. Note that SRN is trained with one-hot encoded primitive tokens as input but operates on token probabilities output by SPN for image-level pre-training. Primitives with incorrect syntax are removed at a post-processing step. We implemented all modules in Pytorch [30] and use the Adam optimizer [20].

### v2

@@ Abstract
This work introduces PICASSO, a framework for the parameterization of 2D CAD sketches from hand-drawn and precise sketch images. PICASSO converts a given CAD sketch image into parametric primitives that can be seamlessly integrated into CAD software. Our framework leverages rendering self-supervision to enable the pre-training of a CAD sketch parameterization network using sketch renderings only, thereby eliminating the need for corresponding CAD parameterization. Thus, we significantly reduce reliance on parameter-level annotations, which are often unavailable, particularly for hand-drawn sketches. The two primary components of PICASSO are (1) a Sketch Parameterization Network (SPN) that predicts a series of parametric primitives from CAD sketch images, and (2) a Sketch Rendering Network (SRN) that renders parametric CAD sketches in a differentiable manner and facilitates the computation of a rendering (image-level) loss for self-supervision. We demonstrate that the proposed PICASSO can achieve reasonable performance even when finetuned with only a small number of parametric CAD sketches. Extensive evaluation on the widely used SketchGraphs [37] and CAD as Language [14] datasets validates the effectiveness of the proposed approach on zero- and few-shot learning scenarios.
@@ 1 Introduction
Computer-Aided Design (CAD) has become the industry norm for mechanical design of any product prior to manufacturing. CAD software [3, 2] enhances the productivity of engineers and enables efficient extension or alteration of existing designs. The modern CAD workflow widely adopts the paradigm of feature-based modeling [44], where initially a series of two-dimensional parametric CAD sketches is specified, followed by CAD operations (e.g. extrusion, revolution, etc.) to form a 3D solid. CAD sketches comprise a collection of geometric primitives (e.g. lines, arcs) as well as constraints enforced between those primitives (e.g. coincident, parallel, etc.). Feature-based CAD constitutes an efficient way of constructing complex 3D models [20, 11, 26, 12]. Commonly, the modelling process starts with conceptualization of a design by roughly drawing it by hand [38]. Designers are tasked with meticulously translating these drawings, often in the form of raster images, into parametric CAD sketches. Depending on the design complexity, the task can be time-consuming even for highly skilled designers. Thus, the automation of the CAD parameterization process (Figure 1-top) has gained attention in the research community [46, 38] and CAD industry [1, 3].
Figure 1: PICASSO is a novel self-supervised framework for CAD sketch parameterization. Unlike methods relying on parametric annotations, PICASSO is pretrained via rendering self-supervision on CAD sketch images only, thus drastically reducing the need for parametrically annotated sketches. We demonstrate PICASSO’s effectiveness in both few-shot and zero-shot settings.
@@
Figure 2: Framework overview. PICASSO is composed of two networks, namely the Sketch Parametrization Network (SPN) and Sketch Rendering Network (SRN). Once trained, SRN is kept frozen and used for rendering self-supervision using a multiscale $l2$ loss. This allows for image-level pre-training of the CAD sketch parameterization network SPN.
@@
Parameterization of CAD sketches from raster images constitutes a complex problem due to the large solution space required to model parametric entities, the nuanced intricacies of sketch designs, and the inherently inaccurate nature of hand-drawings. Compared to vectorized stroke inputs [4], raster images pose additional challenges due to indistinguishable, overlapping or closely positioned primitives on the image domain. The availability of large-scale CAD sketch datasets [37, 14] has enabled tackling this problem with learning-based approaches. Recent works [14, 29, 38] propose autoregressive generative models to learn CAD sketch parameterization through general-purpose language modelling. Similarly to next word prediction for natural language processing (NLP), these models predict the next primitive in the sketch and can be conditioned on images for sketch parameterization. Transformer-based networks are typically trained to predict tokens representing the type and parameters of each primitive via parameter-level supervision. These networks are adept at exploiting large-scale annotated datasets with millions of sketches (e.g. SketchGraphs [37]) (Figure 1-bottom left). Nevertheless, there are real-world situations where this quantity of labeled sketches is unavailable. Examples include rough hand-drawn CAD sketches and 2D cross-sections [6] from 3D scans, where collecting parameter annotations is challenging. These examples exhibit a distribution shift w.r.t. to CAD sketches in large-scale datasets, underscoring the necessity of learning CAD sketch parametrization from raster images when parameter-level annotations are scarce or nonexistent.
@@
In this work, we introduce a framework for Parametric Inference of CAD Sketches via Rendering Self-SupervisiOn, refereed to as PICASSO. PICASSO enables learning parametric CAD sketches directly from precise or hand-drawn images, even when parameter-level annotations are limited or unavailable (Figure 1-bottom right). This is achieved by utilizing the geometric appearance of sketches as a learning signal to pretrain a CAD parameterization network. Specifically, PICASSO is composed of two main components: (1) a Sketch Parameterization Network (SPN) that predicts a set of parametric primitives from CAD sketch images and (2) a Sketch Rendering Network (SRN) to render parametric CAD sketches in a differentiable manner. SRN enables image-to-image loss computation that can be used to pretrain SPN, leading to zero- and few-shot learning scenarios for hand-drawn sketch parametrization (see Figure 1). To the best of our knowledge, we are the first to address CAD sketch parameterization with limited or without parametric annotations. PICASSO can achieve strong parameterization performance with only a small number of annotated samples. Moreover, departing from recent work based on autoregressive modelling [38, 29, 14], the proposed SPN is a feed-forward network. The sketch is treated as an unordered set of primitives and a CAD sequence is predicted in a non-autoregressive manner. Experiments show that SPN outperforms recent autoregressive state-of-the-art on parameterizing precise and hand-drawn CAD sketch images. An overview of PICASSO is presented in Figure 2.
@@ 2 Related Works
Concurrently, SketchGen [29] and Vitruvion [38] propose two-stage architectures for both primitive and constraint generation. Seff et al. [38] also investigate their model’s efficacy for the parameterization of hand-drawn sketches. These methods follow the autoregressive learning strategy as it constitutes the natural choice for generative modelling. Autoregressive inference is also well-suited for the related application of sketch auto-completion but might be suboptimal for CAD parameterization from images (see detailed discussion in Section 4.2). In contrast, the proposed SPN is a feed-forward network. Related to ours is also the non-autoregressive method of [46]. Authors introduce the task of modular concept discovery that is addressed through a program library induction perspective. Recently, the authors in [40] presented a concurrent work to ours. They highlighted the issue of error accumulation in autoregressive models and introduced a feed-forward strategy for CAD sketch parameterization, which is similar to our Sketch Parameterization Network (SPN). Also concurrently to ours, the work of  [19] proposes a single-stage architecture for joint sketch parameterization and constraint prediction. All aforementioned methods solely rely on parameter-level supervision and overlook the rendered geometry of a CAD sketch. In this work, rendering self-supervision is proposed to provide an alternative signal for sketch parameterization. Such signal allows for pre-training CAD sketch parameterization network, hence enabling it to parameterize hand-drawn or precise CAD sketches under limited or without parametric supervision.
@@ 3 Problem Statement
Figure 3: Overview of the Sketch Rendering Network (SRN). SRN is modeled by a transformer encoder-decoder that learns a mapping from parametric primitive tokens to the sketch image domain. SRN enables neural differentiable rendering that we leverage for rendering self-supervision of SPN.
Our goal becomes to learn the mapping $\boldsymbol{F}_{\theta}:\{0,1\}^{h\times w}\rightarrow\mathcal{P}^{n}$ with model parameters $\theta$, from the image domain to the space of parametric primitives. As in [38], each primitive is represented by a set of tokens defining primitive types and quantized primitive parameters. A $6$-bit uniform quantization is adopted. More specifically, every primitive is expressed as a set of 8 tokens $\mathbf{p}_{n}~{}=~{}\{t_{i}^{j}\}_{j\in\llbracket 1..8\rrbracket}$ with $t_{i}^{j}\in\llbracket 0..72\rrbracket$. Tokens in $\llbracket 0..6\rrbracket$ represent primitive types i.e., padding, start, end, arc, circle, line, and point, respectively. Tokens in the range $\llbracket 7..70\rrbracket$ correspond to quantized primitive parameters and values $\{71,72\}$ refer to whether the primitive is a construction primitive (used by designers for referencing). The complete set of quantized primitives can be formally defined as ${\mathbf{y}}~{}\in~{}\{0,1\}^{8n\times 73}$ where $8$ represents the tokens for each primitive and $73$ is the tokenization interval.
@@ 4.1 Sketch Rendering Network
This work proposes a Sketch Rendering Network (SRN) designed for neural rendering of parametric CAD sketches. SRN learns $\mathbf{\Phi}_{\phi}:~{}\mathcal{P}^{n}~{}\rightarrow~{}\{0,1\}^{h\times w}$, i.e., the mapping from a set of parametric primitive tokens to the sketch image domain. The primary utility of SRN is to allow the computation of an image-to-image loss between predicted primitives and their corresponding precise or hand-drawn input sketch image, thus enabling image-level pre-training for CAD sketch parametrization. Note that explicit rendering of CAD sketches implies the direct rasterization of parametric primitives, and it is inherently a non-differentiable process. While differentiable rendering solutions exist in literature [23], their use is limited to parameter refinement.
Network Architecture: The proposed architecture for SRN is depicted in Figure 3. SRN operates on one-hot input tokens ${\mathbf{y}=[\mathbf{y}_{1}\;\mathbf{y}_{2}\;...\;\mathbf{y}_{8n}]}~{}\in~{}\{0,1\}^{8n\times 73}$. Firstly, input tokens are individually projected through a linear layer $\mathbf{\Phi}_{f}$ to produce token embeddings $\mathbf{f}_{r}=[\mathbf{f}_{r_{1}},\mathbf{f}_{r_{2}},...,\mathbf{f}_{r_{8n}}]$, where $\mathbf{f}_{r_{i}}=\mathbf{\Phi}_{f}(\mathbf{y}_{i})~{}\in~{}\mathbb{R}^{d_{q}}$ for the $i$-th token and $d_{q}$ denotes the embedding dimension. The $\mathbf{f}_{r}$ embeddings are concatenated with fixed positional embeddings $\textbf{e}^{pos}_{i}$ [30] and processed by a transformer encoder network $\mathbf{\Phi}_{e}$ without changing their dimensions. In the decoding phase, the decoder $\mathbf{\Phi}_{d}$, inspired by [16], employs self and cross-attention to map a set of randomly initialized learnable patch queries $\mathbf{q}_{\Phi}~{}\in~{}\mathbb{R}^{\frac{hw}{d_{e}^{2}}\times d_{e}\times d_{e}}$ to a set of decoded patch embeddings $\mathbf{z}=[\mathbf{z}_{1}\;\mathbf{z}_{2}\;...\;\mathbf{z}_{\frac{hw}{d_{e}^{2}}}]~{}\in~{}\mathbb{R}^{\frac{hw}{d_{e}^{2}}\times d_{e}\times d_{e}}$. Patch queries $\mathbf{q}_{\Phi}$ serve as learnable tokens that query the encoded representations to extract spatial and structural information that is required for the rendering. Finally, each decoded patch embedding $\mathbf{z}_{i}$ pass through a patch-wise linear layer with a sigmoid activation function, to produce the corresponding output patch of pixel-intensities $\mathbf{o}_{\frac{hw}{d_{e}^{2}}}~{}\in[0,1]^{d_{e}\times d_{e}}$. The final rendered image is reconstructed by assembling the output patches $\mathbf{o}_{i}$.
@@ 4.2 Sketch Parametrization Network
Network Architecture: Figure 4 shows the architecture of SPN. First, a convolutional [35] backbone $\mathbf{\Psi}_{f}$ extracts a set of image features $\mathbf{f}\in\mathbb{R}^{c\times h\times w}$, for input sketch image $\mathbf{X}$. The input feature map is divided into non-overlapping patches that are embedded along with fixed positional embeddings [30] and passed through a standard vision transformer encoder [10] $\mathbf{\Psi}_{v}$ to produce a set patches $\mathbf{v}~{}\in~{}\mathbb{R}^{\frac{hw}{d_{e}^{2}}\times d_{e}\times d_{e}}$. The decoder $\mathbf{\Psi}_{d}$ is a vanilla transformer [39] trained to decode a set of $d$ randomly initialized learnable query vectors $\mathbf{q}_{\Psi}~{}\in~{}~{}\mathbb{R}^{d\times d_{q}}$ into primitive tokens via self and cross-attention. Note that $d_{e}$ and $d_{q}$ maintain the same value as for SRN. The decoded tokens are classified independently via a linear layer with softmax activation resulting in token probabilities $\hat{\mathbf{y}}~{}\in~{}[0,1]^{8n\times 73}$.
@@ 4.3 Rendering Self-Supervision for CAD Sketch Parameterization
$\mathcal{L}_{ml2}=\sum_{s\in S}\|d_{s}\left(\mathbf{\Phi}_{\phi}(\boldsymbol{F}_{\theta}(\mathbf{\hat{X}})\right))-d_{s}\left(\mathbf{X}\right)\|_{2}^{2}\ ,$ (1)
@@ 5 Experiments
Implementation Details: Images are of size $128\times 128$. SPN uses a U-Net [35] convolutional backbone with a ResNet34 [17] that produces a $16\times 128\times 128$ feature map. For SPN, the transformer encoder-decoder is formed by $4$ layers each, with $8$ heads and $d_{q}=256$ latent dimensions. The SRN transformer has $12$ layers with $8$ heads and the same latent dimension as SPN. Patch size $d_{e}$ is set to $16$. As in [38], the maximum number of primitives is fixed to $n=16$. We train SRN and SPN for $40$ and $20$ epochs respectively, with a learning rate of $1\times 10^{-4}$ and a batch size of $128$. Note that SRN is trained with one-hot encoded primitive tokens as input but operates on token probabilities output by SPN for image-level pre-training. Primitives with incorrect syntax are removed. We implemented modules in Pytorch [31] and use the Adam optimizer [21].